OpenStudy (anonymous):

Sketch a triangle with lengths of sides 4, 5, and 6 units. Is the triangle acute, right, or obtuse? How do you know?

Mehek (mehek14):

did you sketch it?

OpenStudy (anonymous):

xD good idea

Mehek (mehek14):

:)

OpenStudy (anonymous):

but how can i use the pythagorean theorem?

Mehek (mehek14):

\(\Large{4^2 + 5^2 = 6^2}\) try it

Mehek (mehek14):

\(\Large{4^2?}\)

OpenStudy (anonymous):

ok so 16 and 25

Mehek (mehek14):

add them

OpenStudy (anonymous):

but what do you do with 41?

OpenStudy (anonymous):

u find the square root right?

Mehek (mehek14):

ok first tell me what is \(\Large{6^2}\)

OpenStudy (anonymous):

36

Mehek (mehek14):

so is \(\Large{41 = 36 ?}\)

OpenStudy (anonymous):

No

Mehek (mehek14):

so not a right triangle

OpenStudy (anonymous):

ok i knew that but how do i find out if it is an obtuse or acute

OpenStudy (anonymous):

if its bigger or smaller than C^2?

Mehek (mehek14):

sketch it first

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

very well

OpenStudy (anonymous):

just wondering if there is an easier way mathematically

OpenStudy (anonymous):

than to have to sketch and change the sketch so all sides are appropriately drawn

OpenStudy (anonymous):

so its obtuse?

Mehek (mehek14):

I found this: Compare the hypotenuse to the longest side given in the problem. If the longest given side is shorter than the hypotenuse, the given triangle is acute. If it is longer than the hypotenuse, the given triangle is obtuse.

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

the hypotenuse is 6?

Mehek (mehek14):

\(\Large{5^2 < 6^2~right?}\)

OpenStudy (anonymous):

yes true

OpenStudy (anonymous):

therefore, it is obtuse

Mehek (mehek14):

the given sides are 4 and 5 the longest is 5 since it is less than the hypotenuse, it is an acute triangle

OpenStudy (anonymous):

4,5, and 6 right?

Mehek (mehek14):

the sides are 4 and 5 the hypotenuse is 6 5<6 so acute

OpenStudy (anonymous):

how did you find the hypotenuse?

OpenStudy (anonymous):

sorry

Mehek (mehek14):

\(\Large{4^2 + 5^2 = 6^2}\)

OpenStudy (anonymous):

but \[4^2+5^2\neq6^2\]

Mehek (mehek14):

\(\Large{5^2 = 25}\) \(\Large{6^2 = 36}\) \(\Large{25<36}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thx

Mehek (mehek14):

since the longest given side is < the hypotenuse, it is an acute triangle

OpenStudy (anonymous):

ok

Mehek (mehek14):

i gtg hope i helped

OpenStudy (anonymous):

yep it did